OpenStudy (anonymous):

Find all real values of λ, if any, for which the system has a nontrivial solution. 2x1 + 3x2 = λx1 4x1 + 3x2 = λx2

OpenStudy (fibonaccichick666):

What do your 'x's mean?

OpenStudy (anonymous):

they are the variables in the system

OpenStudy (anonymous):

\[2x_{1} + 3x _{2} = \lambda x _{1}\]

OpenStudy (fibonaccichick666):

ah ok

OpenStudy (anonymous):

\[4x _{1} + 3x _{2} = \lambda x _{2}\]

OpenStudy (fibonaccichick666):

What can you tell me about how many equations you need in relation to the number of variables you have?

OpenStudy (anonymous):

with two variables, don't you need two equations?

OpenStudy (fibonaccichick666):

yup, how about 3?

OpenStudy (anonymous):

3... ?

OpenStudy (fibonaccichick666):

and 4?

OpenStudy (anonymous):

equations?

OpenStudy (fibonaccichick666):

4 variables?

OpenStudy (anonymous):

would need 4 equations

OpenStudy (fibonaccichick666):

and n variables?

OpenStudy (anonymous):

n equations

OpenStudy (fibonaccichick666):

good, so now, how many variables do you have?

OpenStudy (anonymous):

2

OpenStudy (fibonaccichick666):

now, do you have enough equations?

OpenStudy (anonymous):

yes

OpenStudy (fibonaccichick666):

(since we are saying \(\lambda\) isn't a real variable)

OpenStudy (anonymous):

no, its acting as as scalar isn't it?

OpenStudy (fibonaccichick666):

yes, sort of.

zepdrix (zepdrix):

Hmm I haven't taken much Linear so I'm not sure how much help I can be. But lemme share what I'm thinking. \[\Large\rm 2x_{1} + 3x _{2} = \lambda x _{1}\]\[\Large\rm 4x _{1} + 3x _{2} = \lambda x _{2}\]Subtract the right hand sides so we have all of our x's on one side. Then writing it in matrix form gives us:\[\Large\rm \left[\begin{matrix}2-\lambda & 3 \\ 4 & 3-\lambda\end{matrix}\right]\vec X =\left(\begin{matrix}0 \\ 0\end{matrix}\right)\]And then we can do something with that matrix yes? Taking the determinate gives us some Lambda values... I don't remember what that means though :P Any ideas Kmendall?

OpenStudy (anonymous):

actually I think I may be able to figure something out from there...

OpenStudy (fibonaccichick666):

Oh I was assuming no linear alg. that makes life easier, ok

OpenStudy (anonymous):

sorry, this is a linear algebra question!

OpenStudy (fibonaccichick666):

make the modified variable matrix, then get it down to reduced row echelon and you should see some lambda values

OpenStudy (anonymous):

ok, I'm having a moment... how do I make the modified variable matrix?

OpenStudy (fibonaccichick666):

it's essentially what zep gave ya,

OpenStudy (fibonaccichick666):

just put it all together so it's a 2 by 3

OpenStudy (anonymous):

ok, that's what I was thinking, just doubting myself Thanks

OpenStudy (fibonaccichick666):

you got it